OpenStudy (anonymous):

Let f(x) = 7x - 13. Find f-1(x).

OpenStudy (mertsj):

y=7x-13 Switch the x and the y and solve for y

OpenStudy (hoblos):

y=7x -13 switch x and y x=7y-13 y= (x+13)/7 f^-1(x)=(x+13)/7